Incorrect Question 5 0/1 pts A 25.0 mL sample of 0.723 M HCIO4 is titrated with 0.27 M KOH solution. The H30* concentration aIncorrect Question 6 0/1 pts What could be the pKof this weak acid according to the titration curve below? Weak Acid TitratioPlease explain the correct answer for both questions. Thank you!
Answer #1

Answer 5

Correct answer is b)

Explanation:

Moles of HClO​​​​​4= 0.723 M*25 mL = 18.075 mmol

Moles of KOH = 0.270 M*80 mL = 21.600 mol

As moles of KOH are more, so, excess moles of KOH = (21.600 - 18.075) millimol

= 3.525 millimol

Total volume = 25 ml + 80 ml = 105 ml

[OH-] = 3.525 millimol/105 ml = 0.0336 M

[H+] = 10-14/0.0336 = 29.76 x 10-14 M

= 3.0 x 10-13 M

Answer 6

Correct answer is d) 7

Explanation: pKa is the pH at the middle region of least steep curve.